OpenStudy (anonymous):

If f(x) is even and 1−1(f(x) − 7) dx = 6, find 1 0 f(x) dx.

OpenStudy (ranga):

The question is not clear. You can use the Equation button below the reply box to type math equations.

OpenStudy (anonymous):

\]\[If f(x) is even and \int\limits_{-1}^{1}(f(x) − 7) dx = 6, find \int\limits_{0}^{1} f(x) dx.\]

OpenStudy (ranga):

Split the integral on the left to two parts. And for the limits go from -1 to 0 and 0 +1. I am assuming for even function; f(x) = f(-x). So the -1 to 0 can be folded to 0 to + 1. simplify.

OpenStudy (dan815):

oh - oops

OpenStudy (dan815):